Let T be a linear operator whose matrix with
respect to the basis {(1,1,1),(1,-1,0),(1,0,-1)} is the matrix [
4 0 0
0 1 0
0 0 1. ]
Obtain the matrix of T wrt the standard basis.
1
Expert's answer
2020-12-15T13:33:03-0500

"\begin{bmatrix}
 1& 1& 1 \\
 1& -1& 0 \\
 1& 0& -1 \\
\end{bmatrix};" "\begin{bmatrix}
 4& 0& 0 \\
 0& 1& 0 \\
 0& 0& 1 \\
\end{bmatrix};"


"\begin{bmatrix}
 1& 1& 1 \\
 1& -1& 0 \\
 1& 0& -1 \\
\end{bmatrix} \times T=
\begin{bmatrix}
 4& 0& 0 \\
 0& 1& 0 \\
 0& 0& 1 \\
\end{bmatrix};"


"T=\begin{bmatrix}
 1& 1& 1 \\
 1& -1& 0 \\
 1& 0& -1 \\
\end{bmatrix} ^{-1}
\begin{bmatrix}
 4& 0& 0 \\
 0& 1& 0 \\
 0& 0& 1 \\
\end{bmatrix};"



Consider the matrix "\begin{bmatrix}
 1& 1& 1 \\
 1& -1& 0 \\
 1& 0& -1 \\
\end{bmatrix}" is A. We use Linear Row Reduction to find the Inverse Matrix of A:

"A^{-1}=\begin{bmatrix}
 1& \ \ \ 1&1 \ \ \ \ \ \ \ \ \ \ \ \ \ 1&0&0 \\
 1& -1& 0 \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0&1&0 \\
 1&\ \ \ 0& -1 \ \ \ \ \ \ \ \ \ \ \ 0&0&1 \\
\end{bmatrix} ^{-1}" = "\begin{Bmatrix}
 r_1=\frac{r_1+r_2+r_3}{3} \\
 r_2=\frac{r_1+r_3-2\times r_2}{3} \\
 r_3=\frac{r_1+r_2-2\times r_3}{3}
\end{Bmatrix}" =


"A^{-1}=\begin{bmatrix}
1&0&0\ \ \ \ \ \ \ \ \ \ \frac{1}{3} & \frac{1}{3} & \frac{1}{3} \\
0&1&0\ \ \ \ \ \ \ \ \ \ \frac{1}{3} & -\frac{2}{3} & \frac{1}{3} \\
0&0&1\ \ \ \ \ \ \ \ \ \ \frac{1}{3} & \frac{1}{3} & -\frac{2}{3} \\
\end{bmatrix} ;"


"T=\begin{bmatrix}
 \frac{1}{3} & \frac{1}{3} & \frac{1}{3} \\
 \frac{1}{3} & -\frac{2}{3} & \frac{1}{3} \\
 \frac{1}{3} & \frac{1}{3} & -\frac{2}{3} \\
\end{bmatrix} 
\begin{bmatrix}
 4& 0& 0 \\
 0& 1& 0 \\
 0& 0& 1 \\
\end{bmatrix};"


"T=\begin{bmatrix}
 \frac{4}{3} & \frac{1}{3} & \frac{1}{3} \\
 \frac{4}{3} & -\frac{2}{3} & \frac{1}{3} \\
 \frac{4}{3} & \frac{1}{3} & -\frac{2}{3} \\
\end{bmatrix} ;" .